OpenStudy (anonymous):

write the equasion for the inverse function y=arcsin3x

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(y=\arcsin 3x\) start by solving x :- take \(\sin\) both sides \(\sin y=\sin ~\arcsin 3x\) \(\sin y= 3x\) divide 3 both sides \(\large \frac{\sin y}{3} = x\) swap x and y \(\large \frac{\sin x}{3} = y\)

ganeshie8 (ganeshie8):

so the inverse function is \(\large y = \frac{\sin x}{3} \) make sure u understood the steps, u wil see lot of these in this module